OpenStudy (anonymous):

Thank you! Can you assist me with this too? 2x*3x+2/5z*3z-3

jimthompson5910 (jim_thompson5910):

is this \[\Large \frac{2x*3x+2}{5z*3z-3}\] ??

jimthompson5910 (jim_thompson5910):

yes? no?

OpenStudy (anonymous):

Is this a rational expression? I am not sure if it is correct or not. I do nt see how it can be simplified any further than it is at this point...am I right?

jimthompson5910 (jim_thompson5910):

well you can multiply terms to go from \[\Large \frac{2x*3x+2}{5z*3z-3}\] to \[\Large \frac{6x^2+2}{15z^2-3}\] but that's about it

OpenStudy (anonymous):

THANK YOU FOR HELPING!!

jimthompson5910 (jim_thompson5910):

np